complete the square, and find the roots of the quadratic equation.
x^2 + 16x = 0

Answer by venugopalramana(3286)

complete the square, and find the roots of the quadratic equation.
x^2 + 16x = 0...USE (A+B)^2=A^2+2AB+B^2
X^2+2(X)(8)+8^2-8^2=0
(X+8)^2=8^2
x+8=8...................OR X+8=-8
X=0................OR .....X=-16